# Autoregressive models for panel data causal inference with application to state-level opioid policies

### v1

@@ 7 Discussion
We examined the impact of four policies on opioid prescribing, highlighting meaningful evidence that must access PDMPs decreases opioid-prescribing rates within a state. We also found null effects showing that naloxone access laws and medical marijuana did not meaningfully reduce opioid prescriptions at the state level, and that medical marijuana may reduce the impact of must access PDMPs when enacted jointly. Taken together, our findings suggest that a combination of policies may be needed to effectively address the opioid crisis. In the end, the application must be interpreted with caution when considering some of the null findings for the other policies. The policies we examine were enacted to improve the lives of individuals in the United States, and even small effects of policies could translate to saving or helping thousands of individuals.
@@
Our work also makes a strong methodological contribution. In particular, we propose a novel causal framework that provides formal justification for using debiased autoregressive models to estimate causal effects. While our causal assumptions make strong restrictions on effect heterogeneity, we show mathematically that the presence of effect heterogeneity likely only causes small bias on our effect estimates. Moreover, we show how our approach easily extends to allow for identification and estimation in the presence of multiple policy changes. By contrast, effect estimation in this context is more challenging for difference-in-differences and synthetic controls methods, as discussed above in Section 4.1. Finally, we provide additional simulation evidence showing that our proposed method outperforms existing methods in opioid-policy settings where we have small sample sizes with a changing policy landscape, consistent with previous studies (Griffin et al., 2021; Schell, Griffin and Morral, 2018; Griffin et al., 2023, 2022). More generally, the proposed debiased autoregressive models offer a promising method for policy researchers to evaluate the effects of social and health policies in multiple applied areas, including gun violence, opioids, and COVID-19 policy evaluations. By providing relatively low variance estimates in small data settings, this method allows for more precise causal inferences with relatively low bias.
@@
On the other hand, this method is not without limitations. Our identification and estimation strategy requires making strong functional form assumptions on the treatment effect function itself, a problem that may be difficult to reason about in certain applications. Additionally, while we have shown that effect heterogeneity is likely only to induce small bias with respect to effect estimation in a highly simplified setting, we have not formally shown that these results extend in more complex settings where we condition on multiple pre-treatment outcomes, covariates, and attempt to evaluate the effects of multiple policies. Valuable future research may therefore consider how to conduct formal sensitivity analyses with respect to the required causal assumptions, and to formalize how heterogeneity may affect model-based estimates in more complex settings. Nevertheless, we believe that we have made strong progress in formally justifying the use of debiased autoregressive models for causal effect estimation and for further arguing for their utility in applied policy settings.
@@ Acknowledgements
This research was financially supported through a National Institutes of Health (NIH) grant (P50DA046351) to RAND (PI: B.D.S.) as well as Arnold Ventures (AV). NIH and AV had no role in the design of the study, analysis, and interpretation of data nor in writing the manuscript. We wish to thank Andrew Morral, Megan Schuler, Bradley Stein, Rosalie Pacula, David Powell, and Evan Peet for comments on different iterations of this paper and/or the execution of the policy evaluation. We also thank the members of the OPTIC Advisory Board for their feedback on these analyses.
@@ Appendix A Causal assumptions for multiple policies
We define our target parameter as the counterfactual quantity $\mathbb{E}[Y_{it}(\bar{a},\bar{m})]$. To identify this quantity in terms of the observed datam assume the following:
@@
$\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m}\implies Y_{it}=Y_{it}(\bar{a},\bar{m}),\qquad\forall\bar{a},\bar{m}$
This assumption extends the previous assumption to account for multiple policies. Similar to before, this assumption precludes interference across states (or units) – that is, each states’ observed outcome is only a function of the policies implemented in that state.
@@
This again extends the previous assumption to incorporate multiple policies. The key difference is that carryover effects are limited to the previous $l$ time-periods for both policies.
@@
The multiple policy ignorability assumption simply states that both policies are jointly randomized with respect to the potential outcomes given the potential outcomes absent treatment. Note that this assumption allows for causal pathways between the two policies, so that $M$ ($A$) may lie on the pathway from $A$ ($M$) to $Y$.
@@
$Y_{it}(\bar{a},\bar{m})=Y_{it}(0,0)+\Delta_{t}(\bar{a},\bar{m})$
This assumption again implies that there is no effect heterogeneity on the additive scale (though this could be modified to the multiplicative scale if desired).
@@
For some $\epsilon>0$ and $e(y)=P\left\{\bar{A}_{it},\bar{M}_{it}\mid Y_{i,t-1}(\bar{0})=y\right\}$, we have
$P[\min_{\bar{a},\bar{m}}P\left[e\{Y_{i,t-1}(0)\}\geq\epsilon\right]]=1.$
This assumption again simply extends the previous positivity assumption by stating that there is some positive probability of observing any treatment history value for either treatment conditional on a specific value of the previous potential outcome absent any treatments. Under these assumptions, we then obtain the identification result as follows:
$\Delta_{t}$ $=\mathbb{E}[\mathbb{E}[Y_{it}\mid\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m},Y_{i,t-1}(0)]-\mathbb{E}[Y_{it}\mid\bar{A}_{it}=\bar{0},\bar{M}_{it}=\bar{0},Y_{i,t-1}(0)]]$
$=\mathbb{E}[\mathbb{E}(Y_{i,t}\mid\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m},Y_{i,t-1}-\Delta_{t-1}(\bar{A}_{t-1},\bar{M}_{t-1}))$
$-\mathbb{E}(Y_{i,t}\mid\bar{A}_{it}=\bar{0},\bar{M}_{it}=\bar{0},Y_{i,t-1}-\Delta_{t-1}(\bar{A}_{t-1},\bar{M}_{t-1}))].$
@@ Appendix B More details on differential measurement error
In this section, we allow the measurement error to depend on any of the terms in the model, and we highlight the bias obtained in this more general scenario than what was considered in the manuscript. This results in a more general expression of the bias compared to the bias expression seen in the manuscript. Specifically, one can show that our estimate of the treatment effect in the presence of measurement error induced by treatment effect heterogeneity converges to
@@
If the measurement error is classical, then any covariance terms involving $U_{i,t-1}$ are zero and therefore the measurement error only impacts the $\text{var}(Y_{i,t-1}(0)+U_{i,t-1})$ term, which simplifies the formula to what was seen in equation (4) of the main manuscript. If, however, the measurement error is differential, the bias can take different magnitudes or directions. Differential measurement error here could occur if $U_{i,t-1}$ is correlated with any of $A_{it},Y_{it},$ or $Y_{i,t-1}(0)$. Therefore it is worth considering in our context whether we expect the bias to be differential, if we have any prior knowledge about the direction of bias, and whether it depends on the source of heterogeneity.
@@
The last form of differential measurement error to consider is the case where $U_{i,t-1}$ is correlated with $Y_{i,t-1}(0)$. This could occur if units that have larger values of $Y_{it}(0)$ are more likely to benefit from treatment. In the opioid setting, states with higher mortality from opioid use might be more likely to benefit from polices reducing its impact. One thing to note is that dependence between $U_{i,t-1}$ and $Y_{i,t-1}(0)$ makes it very likely to have dependence between $U_{i,t-1}$ and $Y_{it}$ as well given the clear relationship between potential outcomes and observed outcomes. For this reason, many of the issues mentioned in the manuscript for dependence with $Y_{it}$ would also apply in this situation; however, here we focus specifically on the impacts of the dependence between $U_{i,t-1}$ and $Y_{i,t-1}(0)$. Letting $P_{y}(U_{i,t-1})$ denote the projection of $U_{i,t-1}$ onto $Y_{i,t-1}(0)$, the variable we adjust for can be written as
@@ Appendix C Inference on sample level estimands
These approaches differ, however, in defining the posterior distribution for $Y_{it}^{*}(a)$. The first approach sets $Y_{it}^{*}(A_{it})=Y_{it}$ and draws $Y_{it}^{*}(1-A_{it})$ from an appropriate distribution (for example, normal or negative binomial) with parameters given by our model. This approach acknowledges that one of the counterfactuals is observed and does not need to be imputed, and then imputes the other from the model. The second approach draws both $Y_{it}^{*}(1)$ and $Y_{it}^{*}(0)$ from the posterior predictive distribution; for example, a normal or negative binomial distribution with parameters given by our model. This approach ignores information from the observed counterfactual and instead imputes both assuming they are independent. The first approach may seem more natural as it uses the observed potential outcome. In our experience, however, this approach works well when the model is correctly specified, but suffers more from moderate amounts of model misspecification relative to the second approach. Many components of the model may be misspecified, including the treatment effect function. If components other than the treatment effect portion of the model are misspecified, we will obtain poor predictions of $Y_{it}^{*}(1-A_{it})$, in turn leading to poor estimates of $SATT_{t}$. The second approach instead uses the difference in two posterior predictive draws, $Y_{it}^{*}(1)-Y_{it}^{*}(0)$: this difference may better estimate the treatment effect, even when individually each of the draws $Y_{it}^{*}(a)$ are biased for the true potential outcome. While both approaches are reliant to some degree on correct model specification, the first approach relies on estimating the full data generating process well, while the second is slightly less reliant on the entire model, but more on the treatment effect specification.
@@ D.1 Data generation mechanism for simulation study
Here, we describe in more detail the specifics of the simulation study used in the manuscript. To vary the degree of confounding, we assign treatment to units with higher values of the outcome, and allow treatment times to vary as a function of the outcome as well. We first calculate the outcome for each unit in time period 7, and then used max-min scaling to let all these values fall between 0 and 1. We refer to this value as $\widetilde{Y}_{7}$. To generate start times that better reflects staggered adoption, we randomly sample $20$ numbers between 8 and 14 with replacement and sort them in increasing order. We can refer to this ordered vector of potential times as $\widetilde{T}_{1},\dots,\widetilde{T}_{20}$ where we have that $\widetilde{T}_{1}\leq\widetilde{T}_{2}\leq\dots\leq\widetilde{T}_{20}$. We assign the $j^{th}$ time $\widetilde{T}_{j}$ to be the start time for unit $i$ with probability proportional to $\widetilde{Y}_{7,i}^{c}$, where $c$ is what we refer to as the confounding constant. Once a unit has been assigned a start time, they are removed from the pool of units moving forward and we continue until 20 units have a treatment time. This process ensures that units with higher values of the outcome are more likely to initiate treatment earlier, and the units that are treated have higher outcome values generally. When $c=0$ there is no confounding by $Y_{it}(0)$ and units and times are chosen completely randomly, but as $c$ increases, the degree of confounding increases. To facilitate comparison, we also ensure that units remain treated for all time periods after treatment initiation, though this is not required for our approach.
@@
$R_{it}^{*}=R_{it}+A_{it}(\tau+\tau_{i}).$
Throughout, we set $\tau=-0.02$ to fix a small average impact of the policy on rates of prescribing, and we vary $\tau_{i}$ depending on what level of heterogeneity we are considering. We set $\tau_{i}=k\boldsymbol{X}_{i}\boldsymbol{\gamma}$ where $\boldsymbol{X}_{i}$ are time-invariant characteristics, and the elements of $\boldsymbol{\gamma}$ are generated from independent standard normal distributions. We explore $k\in\{0,0.009,0.026\}$, which we denote by no heterogeneity, moderate heterogeneity, and high heterogeneity. The two non-zero values were chosen to ensure that $P(|\tau_{i}|>|\tau|)$ was equal to 0.1 or 0.33, respectively. This implies that in the high heterogeneity setting, approximately one third of the subjects have unit-specific treatment effects that are either of the opposite sign of the average treatment effect, or are at least twice as large. The observed outcomes are set as $Y_{it}^{*}=R_{it}^{*}P_{it}$, where $P_{it}$ is the population size for unit $i$ at time $t$, and outcomes are rounded to the nearest integer to ensure count outcomes.
@@ D.2 Negative binomial simulations on opioid prescribing
Here we evaluate the performance of the proposed autoregressive approach when using negative binomial models. As in the simulation section of the manuscript, we utilize the observed opioid data to perform simulation studies instead of using fully synthetic data. We follow the same simulation structure as in Section 5 of the manuscript that allows for staggered adoption of the policy and confounding by previous time points. The only difference between the data generation in the manuscript and in this section is how the treatment effects are generated. As in the manuscript, a small average treatment effect is incorporated, and we vary the degree of heterogeneity of the treatment effect across units. We choose the degree of heterogeneity in the same manner as in the manuscript, but we generate new rates on the multiplicative scale via:
$\frac{R_{it}^{*}}{R_{it}}=\exp\big{(}A_{it}(\tau+\tau_{i})\big{)},$
so that Assumption 9 is satisfied when there is no heterogeneity ($\tau_{i}=0$ for all $i$). As in the manuscript, the constant treatment effects assumption will be violated in the moderate and high heterogeneity situations. We target the same estimand as in Section 5, given by
$\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}\sum_{m=0}^{2}\bigg{\{}Y_{i,T_{i}+m}(1)-Y_{i,T_{i}+m}(0)\bigg{\}}.$
@@
The results can be found in Figures 10 - 13. Overall, the results mirror those seen in Section 5 when using the linear autoregressive models rather than the negative binomial autoregressive models. The improved efficiency leads to large decreases in MSE compared with existing estimators. The synthetic control estimator performs slightly better in terms of MSE in some of the high heterogeneity simulations, though the autoregressive model has the lowest MSE in all other scenarios. The interval coverage for the autoregressive model is near the nominal rate, except in situations with very large degrees of heterogeneity, which echoes the same findings from Section 5 of the manuscript. Lastly, the autoregressive approach has the highest power of all four of the approaches to identify a treatment effect across all simulation scenarios, due to the improved efficiency of the approach.

### v2

@@ 8 Discussion
We examined the impact of four policies on opioid prescribing, highlighting meaningful evidence that must access PDMPs decreases opioid-prescribing rates within a state. We also found null effects showing that naloxone access laws and medical marijuana did not meaningfully reduce opioid prescriptions at the state level, and that medical marijuana may reduce the impact of must access PDMPs when enacted jointly. Taken together, our findings suggest that a combination of policies may be needed to effectively address the opioid crisis. In the end, the application must be interpreted with caution when considering the null findings for the other policies. The policies we examine were enacted to improve the lives of individuals in the United States, and even small effects of policies could translate to saving or helping thousands of individuals.
@@
Our work also makes a strong methodological contribution. In particular, we propose a novel causal framework that provides formal justification for using debiased autoregressive models to estimate causal effects. While our causal assumptions make strong restrictions on effect heterogeneity, we show mathematically that the presence of effect heterogeneity likely only causes a small bias. Moreover, we show how our approach can allow for identification and estimation in the presence of multiple policy changes. By contrast, effect estimation in this context is more challenging for difference-in-differences and synthetic controls methods. Finally, we provide simulation evidence showing that our proposed method outperforms existing methods in opioid-policy settings where we have small sample sizes with a changing policy landscape, consistent with previous studies (Griffin et al., 2021; Schell, Griffin and Morral, 2018; Griffin et al., 2023, 2022). More generally, the proposed debiased autoregressive models offer a promising method for policy researchers to evaluate the effects of social and health policies in multiple applied areas, including gun violence, opioids, and COVID-19 policy evaluations. By providing relatively low variance effect estimates in small data settings, this method allows for more precise causal inferences with relatively low bias.
@@
On the other hand, this method is not without limitations. Our identification and estimation strategy requires making strong functional form assumptions on the treatment effect function itself, a problem that may be difficult to reason about in certain applications. Additionally, while we have shown that effect heterogeneity is likely only to induce small bias with respect to effect estimation in a highly simplified setting, we have not formally shown that these results extend in more complex settings where we condition on multiple pre-treatment outcomes, covariates, and where we attempt to evaluate the effects of multiple policies. Valuable future research may therefore consider how to conduct formal sensitivity analyses with respect to the required causal assumptions, and to formalize how heterogeneity may affect model-based estimates in more complex settings. Nevertheless, our results make strong progress in formally justifying the use of debiased autoregressive models for causal effect estimation and for their utility in applied policy settings.
@@ Acknowledgements
This research was financially supported through a National Institutes of Health (NIH) grant (P50DA046351) to RAND (PI: B.D.S.) as well as Arnold Ventures (AV). NIH and AV had no role in the design of the study, analysis, and interpretation of data nor in writing the manuscript. We wish to thank Andrew Morral, Megan Schuler, Bradley Stein, Rosalie Pacula, David Powell, Mark Sorbero, and Evan Peet for comments on different iterations of this paper and/or the execution of the policy evaluation. We also thank the members of the OPTIC Advisory Board for their feedback on these analyses.
@@ Appendix A Causal assumptions for multiple policies
We define our target parameter as the counterfactual quantity $\mathbb{E}[Y_{it}(\bar{a},\bar{m})]-\mathbb{E}[Y_{it}(\bar{a}^{\prime},\bar{m}^{\prime})]$. To identify this quantity in terms of the observed data, we first assume consistency,
@@
$\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m}\implies Y_{it}=Y_{it}(\bar{a},\bar{m}),\qquad\forall\bar{a},\bar{m}.$
This assumption extends the previous assumption to account for multiple policies. Similar to before, this assumption precludes interference across states (or units) – that is, each states’ observed outcome is only a function of the policies implemented in that state. We next assume $l$-limited carryover effects.
@@
This extends the previous l-limited carryover effects assumption to incorporate multiple policies. The key difference is that carryover effects are limited to the previous $l$ time-periods for both policies.
@@
The multiple policy ignorability assumption states that both policies are jointly randomized with respect to the potential outcomes given the potential outcomes absent treatment, and allows for causal pathways between the two policies, so that $M$ (AA) may lie on the pathway from AA ($M$) to $Y$. We next assume constant treatment effects,
@@
$Y_{it}(\bar{a},\bar{m})=Y_{it}(0,0)+\theta_{t}(\bar{a},\bar{m})$
This assumption again precludes effect heterogeneity on the additive scale (though this could be modified to the multiplicative scale if desired).
@@
For some $\epsilon>0$ and $e(y,\bar{a},\bar{m})=\mathbb{P}\left\{\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m}\mid Y_{i,t-1}(\bar{0})=y\right\}$, we have
$\min_{\bar{a},\bar{m}}P\left[e\{Y_{i,t-1}(0),\bar{a},\bar{m}\}\geq\epsilon\right]=1.$
This assumption again simply extends the previous positivity assumption by stating that there is some positive probability of observing any treatment history value for either treatment conditional on a specific value of the previous potential outcome absent any treatments. Finally, we extend assumption 6:
Assumption 15 (Parametric model).
Assume that for each unit $i$ and time period $t$,
$Y_{it}(\bar{0},\bar{0})=\alpha_{t}+\beta_{t}^{\top}\phi(Y_{i,t-1}(\bar{0},\bar{0}))+\epsilon_{it}.$ (7)
for some parameter vector $(\alpha_{1},\beta_{1}^{\top},\dots,\alpha_{T},\beta_{T})^{\top}\in\mathbb{R}^{T(q+1)}$ and some known function $\phi(x):\mathbb{R}\to\mathbb{R}^{q}$. Let $Z_{it}=[1,A_{it},\phi^{\top}(Y_{i,t-1}(\bar{0},\bar{0}))]$ and assume that $\mathbb{E}[Z_{it}^{\top}Z_{it}]$ has full column rank.
Proposition 4.
Under assumptions 10-15, $\mathbb{E}[Y_{it}(\bar{a},\bar{m})]-\mathbb{E}[Y_{it}(\bar{a}^{\prime},\bar{m}^{\prime})]$ is identified in the data as,
$\theta_{a}^{\star\top}(\bar{a}-\bar{a}^{\prime})+\theta_{m}^{\star\top}(\bar{m}-\bar{m}^{\prime}),$
where $(\theta_{a}^{\star},\theta_{m}^{\star})$ are minimizers of $\mathbb{E}[(Y_{it}-\beta_{t}-\theta_{a}^{\top}\bar{A}-\theta_{m}^{\top}\bar{M}-\beta_{t}^{\top}\phi(Y_{it-1}-\theta_{a}^{\top}\bar{A}_{i,t-1}-\theta_{m}^{\top}\bar{M}_{i,t-1})^{2}]$.
@@ Appendix B More details on differential measurement error
In this section, we allow the measurement error to depend on any of the terms in the model, and highlight the bias obtained in this more general scenario than what was considered in the manuscript. This results in a more general expression of the bias compared to the bias expression seen in Section 4. Specifically, we can show that our estimate of the treatment effect in the presence of measurement error induced by treatment effect heterogeneity converges to
@@
If the measurement error is classical, then any covariance terms involving $U_{i,t-1}$ are zero and therefore the measurement error only impacts the $\text{var}(Y_{i,t-1}(0)+U_{i,t-1})$ term, which simplifies the formula to what was seen in equation (5) of the main manuscript. If, however, the measurement error is differential, the bias can take different magnitudes or directions. Differential measurement error here could occur if $U_{i,t-1}$ is correlated with any of $A_{it},Y_{it},$ or $Y_{i,t-1}(0)$. It is thus worth considering in our context whether we expect the bias to be differential, if we have any prior knowledge about the direction of bias, and whether it depends on the source of heterogeneity.
@@
The last form of differential measurement error to consider is the case where $U_{i,t-1}$ is correlated with $Y_{i,t-1}(0)$. This could occur if units that have larger values of $Y_{it}(0)$ are more likely to benefit from treatment. In the opioid setting, states with higher mortality from opioid use might be more likely to benefit from policies reducing its impact. One thing to note is that dependence between $U_{i,t-1}$ and $Y_{i,t-1}(0)$ makes it very likely to have dependence between $U_{i,t-1}$ and $Y_{it}$ as well given the clear relationship between potential outcomes and observed outcomes. For this reason, many of the issues mentioned in the manuscript for dependence with $Y_{it}$ would also apply in this situation; however, here we focus specifically on the impacts of the dependence between $U_{i,t-1}$ and $Y_{i,t-1}(0)$. Letting $P_{y}(U_{i,t-1})$ denote the projection of $U_{i,t-1}$ onto $Y_{i,t-1}(0)$, the variable we adjust for can be written as
@@ Appendix C Inference on sample level estimands
These approaches differ, however, in defining the posterior distribution for $Y_{it}^{*}(a)$. The first approach sets $Y_{it}^{*}(A_{it})=Y_{it}$ and draws $Y_{it}^{*}(1-A_{it})$ from an appropriate distribution (for example, normal or negative binomial) with parameters given by our model. This approach acknowledges that one of the counterfactuals is observed and does not need to be imputed, and then imputes the other from the model. The second approach draws both $Y_{it}^{*}(1)$ and $Y_{it}^{*}(0)$ from the posterior predictive distribution; for example, a normal or negative binomial distribution with parameters given by our model. This approach ignores information from the observed counterfactual and instead imputes both assuming they are independent. The first approach may seem more natural as it uses the observed potential outcome. In our experience, however, this approach works well when the model is correctly specified, but suffers more from moderate amounts of model misspecification relative to the second approach. Many components of the model may be misspecified, including the treatment effect function. If components other than the treatment effect portion of the model are misspecified, we will obtain poor predictions of $Y_{it}^{*}(1-A_{it})$, in turn leading to poor estimates of $SATT_{t}$. The second approach instead uses the difference in two posterior predictive draws, $Y_{it}^{*}(1)-Y_{it}^{*}(0)$: this difference may better estimate the treatment effect, even when individually each of the draws $Y_{it}^{*}(a)$ are biased for the true potential outcome. While both approaches are reliant to some degree on correct model specification, the first approach relies on estimating the full data generating process well, while the second is slightly less reliant on the entire model, relying more on the treatment effect specification itself.
@@ D.1 Data generation mechanism for simulation study
We describe in more detail the specifics of the simulation study used in the manuscript. To vary the degree of confounding, we assign treatment to units with higher values of the outcome, and allow treatment times to vary as a function of the outcome as well. We first calculate the outcome for each unit in time period 7, and then used max-min scaling to let all these values fall between 0 and 1. We refer to this value as $\widetilde{Y}_{7}$. To generate start times that better reflects staggered adoption, we randomly sample $20$ numbers between 8 and 14 with replacement and sort them in increasing order. We can refer to this ordered vector of potential times as $\widetilde{T}_{1},\dots,\widetilde{T}_{20}$ where we have that $\widetilde{T}_{1}\leq\widetilde{T}_{2}\leq\dots\leq\widetilde{T}_{20}$. We assign the $j^{th}$ time $\widetilde{T}_{j}$ to be the start time for unit $i$ with probability proportional to $\widetilde{Y}_{7,i}^{c}$, where $c$ is what we refer to as the confounding constant. Once a unit has been assigned a start time, they are removed from the pool of units moving forward and we continue until 20 units have a treatment time. This process ensures that units with higher values of the outcome are more likely to initiate treatment earlier, and the units that are treated have higher outcome values generally. When $c=0$ there is no confounding by $Y_{it}(0)$ and units and times are chosen completely randomly, but as $c$ increases, the degree of confounding increases. To facilitate comparison, we also ensure that units remain treated for all time periods after treatment initiation, though this is not required for our approach.
@@
$R_{it}^{*}=R_{it}+A_{it}(\theta+\theta_{i}).$
Throughout, we set $\theta=-0.02$ to fix a small average impact of the policy on rates of prescribing, and we vary $\theta_{i}$ depending on what level of heterogeneity we are considering. We set $\theta_{i}=k\bm{X}_{i}\bm{\gamma}$ where $\bm{X}_{i}$ are time-invariant characteristics, and the elements of $\bm{\gamma}$ are generated from independent standard normal distributions. We explore $k\in\{0,0.009,0.026\}$, which we denote by no heterogeneity, moderate heterogeneity, and high heterogeneity. The two non-zero values were chosen to ensure that $P(|\theta_{i}|>|\theta|)$ was equal to 0.1 or 0.33, respectively. This implies that in the high heterogeneity setting, approximately one third of the subjects have unit-specific treatment effects that are either of the opposite sign of the average treatment effect, or are at least twice as large. The observed outcomes are set as $Y_{it}^{*}=R_{it}^{*}P_{it}$, where $P_{it}$ is the population size for unit $i$ at time $t$, and outcomes are rounded to the nearest integer to ensure count outcomes.
@@ D.2 Negative binomial simulations on opioid prescribing
We evaluate the performance of the proposed autoregressive approach when using negative binomial models. As in the simulation section of the manuscript, we utilize the observed opioid data to perform simulation studies instead of using fully synthetic data. We follow the same simulation structure as in Section 6 of the manuscript that allows for staggered adoption of the policy and confounding by previous time points. The only difference between the data generation in the manuscript and in this section is how the treatment effects are generated. As in the manuscript, a small average treatment effect is incorporated, and we vary the degree of heterogeneity of the treatment effect across units. We choose the degree of heterogeneity in the same manner as in the manuscript, but we generate new rates on the multiplicative scale via:
$\frac{R_{it}^{*}}{R_{it}}=\exp\big(A_{it}(\theta+\theta_{i})\big),$
so that Assumption 9 is satisfied when there is no heterogeneity ($\theta_{i}=0$ for all $i$). As in the manuscript, the constant treatment effects assumption will be violated in the moderate and high heterogeneity situations. We target the same estimand as in Section 6, given by
$\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}\sum_{m=0}^{2}\bigg\{Y_{i,T_{i}+m}(1)-Y_{i,T_{i}+m}(0)\bigg\},$
@@
The results can be found in Figures 12 - 15. Overall, the results mirror those seen in Section 6 when using the linear autoregressive models rather than the negative binomial autoregressive models. The improved efficiency leads to large decreases in MSE compared with existing estimators. The synthetic control estimator performs slightly better in terms of MSE in some of the high heterogeneity simulations, though the autoregressive model has the lowest MSE in all other scenarios. The interval coverage for the autoregressive model is near the nominal rate, except in situations with very large degrees of heterogeneity, which echoes the same findings from Section 6 of the manuscript. Lastly, the autoregressive approach has the highest power of all four of the approaches to identify a treatment effect across all simulation scenarios, due to the improved efficiency of the approach.